OpenStudy (christos):

how would you integrate x^3 / sqrt(1 - x^2 )

OpenStudy (amtran_bus):

I think everyone is asleep.

OpenStudy (l):

This is a non-elementary integral. D:

OpenStudy (christos):

I have found sqrt(1 - x^3 - sqrt(1 - x^2 )/ 3 not sure if its correct ,I think there was a way to check your answer using wolfram or something

OpenStudy (l):

interesting. try differentiating it maybe?

OpenStudy (l):

as far as I know, this shouldn't have a simple form.

OpenStudy (karim728):

what methods have you learned so far

OpenStudy (karim728):

you there?

OpenStudy (christos):

ye

OpenStudy (sshayer):

\[I=\int\limits \frac{ x^3 }{ \sqrt{1-x^2} }dx\] put \[\sqrt{1-x^2}=t,x^2=1-t^2,2x dx=-2t dt\] \[=\int\limits \frac{ x^2 x dx }{ \sqrt{1-x^2} }=\int\limits \frac{ \left( 1-t^2 \right)\left( -t dt \right) }{ t }=-\int\limits \left( 1-t^2 \right)dt\] =?

OpenStudy (sshayer):

i think now you can complete.

OpenStudy (sshayer):

you got it or any doubt.

OpenStudy (karim728):